OpenStudy (anonymous):

Marco is draining his two aquariums. The tanks are the same size. One tank has 7 in. of water in it and is being drained at a rate of 1 in./min. The other tank has 5 in. of water in it and is being drained at a rate of 0.5 in./min. After how many minutes will the tanks contain the same amount of water? Remember to enter a numerical answer.

OpenStudy (anonymous):

7 - t = 5 - 0.5t 2 = 0.5t t = 4 minutes